# Deep RL Dual Sourcing Inventory Management with Supply and Capacity Risk Awareness

### v2

@@ 1 Introduction
We organize rest of the paper as summarized next. In Section 2, we mathematically formulate the dual-sourcing inventory management problem and thereby describe our dual sourcing RL methodology as a solution for the problem. In Section 3, we present our main results from numerical experiments. Additionally, in Section 4, we extend our dual sourcing RL baseline model by introducing a capacity mechanism control strategy and provide evaluation results. Finally, we conclude the paper with discussion on future work in Section 5.
@@ 2.4 The Deep RL Algorithm for Training the Buy Policy
Firstly, recall that parameters $\theta\in\Theta$ at any time $t$ essentially dictates the RL agent’s policy as expressed in eq. (3), therefore our problem reduces to learning optimal parameters $\theta^{\star}\in\Theta$. Specifically, we leverage a Deep Neural Network (DNN) architecture for $\pi(\cdot,\cdot;\theta)$. So, $\theta$ in our DRL framework essentially implies the weights and parameters of the constituent neurons in the policy network. We applied the same DNN architecture as used in Madeka et al. (2022). The detailed training algorithm can be found in Appendix B.
@@ 3.1 Training/Evaluation Configurations
We compare our Dual Sourcing RL (DualSrc-RL) buy policy against several baselines, i.e. BaseStockHorizonTip (BSHT), Tailored Base Surge (TBS), Just-in-time RL (JIT-RL). The first two are classic operation research baselines and the JIT-RL is the RL baseline where the RL model is trained as a single-sourcing model (Madeka et al., 2022). More detailed description of those baselines are reported in Appendix A.1.
@@ 3.2 Main Results and Analytics
From the results above, we can observe that overall dual sourcing strategies DualSrc-RL, TBS are favorable in terms of rewards. Furthermore, DualSrc-RL is most profitable in all the run scenarios, and has $4\%$ reward gains over TBS.
@@ 5 Conclusion and Future Work
Appendix A Related work
Reinforcement Learning for Inventory Control: Recently, the progress of Reinforcement Learning (RL) on various optimization problems (Bello et al., 2016; Silver et al., 2016; Nazari et al., 2018; Liu et al., 2022; Liu, 2022) motivates its application in inventory management problems. In this context, Deep RL approaches have emerged more recently over its other Model-Based counterparts, due to improved computational scalability and generalization performance of Deep Neural Network (DNN) architectures. Furthermore, DRL methods have been shown to achieve performance gains over benchmarks for the multi-period newsvendor problem under fairly realistic assumptions on costs, prices, demand and stationarity (Balaji et al., 2019). It is worth highlighting that DRL for single sourcing problem has been comprehensively investigated with extensive empirical evidences at Amazon (Madeka et al., 2022).
A.1 Baseline Methods
A.1.1 Improved Tailored Base Surge Policy
@@
A.1.2 Other Baseline Policies
@@
Appendix B Training Algorithms
B.1 Training Algorithm for the Buy Policy
@@
B.2 Training Algorithm for the Neural Coordinator
@@
Appendix C Input Features
C.1 Featurization for Buying Policy
  1. 1.
    The current inventory level
  2. 2.
    Previous actions aiu that have been taken
  3. 3.
    Time series features
    (a) Historical availability corrected demand
    (b) Distance to public holidays
    (c) Historical website glance views data
  4. 4.
    Static product features
    (a) Product group
    (b) Text-based features from the product description
    (c) Brand
    (d) Volume
  5. 5.
    Economics of the product - (price, cost etc.)
  6. 6.
    Capacity costs – past costs and forecasted future costs
C.2 Featurization for Neural Coordinator
  1. 1.
    (a) Order quantities
    (b) Inventory
    (c) Availability corrected demand
    (d) Inbound
    (e) All the above, but weighted by inbound and storage volumes
  2. 2.
    Forecasted quantities for next L weeks.
    (a) Mean demand at lead time
    (b) Inventory after expected drain at lead time
    (c) All the above, but weighted by inbound and storage volumes
  3. 3.
    Other time series features
    (a) Distance to public holidays
  4. 4.
    Capacity costs (past costs and forecasted future costs)

### v3

@@ 1 Introduction
We organize rest of the paper as summarized next. Section 2 reviews the related work. In Section 3, we mathematically formulate the dual-sourcing inventory management problem and thereby describe our dual sourcing RL methodology as a solution for the problem. In Section 4, we present our main results from numerical experiments. Additionally, in Section 5, we extend our dual sourcing RL baseline model by introducing a capacity mechanism control strategy and provide evaluation results. Finally, we conclude the paper with discussion on future work in Section 6.
2 Related work
Reinforcement Learning for Inventory Control: Recently, the progress of Reinforcement Learning (RL) on various optimization problems (Bello et al., 2016; Silver et al., 2016; Nazari et al., 2018; Liu et al., 2022; Liu, 2022; Liang et al., 2024; Shao et al., 2024) motivates its application in inventory management problems. In this context, Deep RL approaches have emerged more recently over its other Model-Based counterparts, due to improved computational scalability and generalization performance of Deep Neural Network (DNN) architectures. Furthermore, DRL methods have been shown to achieve performance gains over benchmarks for the multi-period newsvendor problem under fairly realistic assumptions on costs, prices, demand and stationarity (Balaji et al., 2019). It is worth highlighting that DRL for single sourcing problem has been comprehensively investigated with extensive empirical evidences at Amazon (Madeka et al., 2022).
@@ 3.4 The Deep RL Algorithm for Training the Buy Policy
Firstly, recall that parameters $\theta\in\Theta$ at any time $t$ essentially dictates the RL agent’s policy as expressed in eq. (3), therefore our problem reduces to learning optimal parameters $\theta^{\star}\in\Theta$. Specifically, we leverage a Deep Neural Network (DNN) architecture for $\pi(\cdot,\cdot;\theta)$. So, $\theta$ in our DRL framework essentially implies the weights and parameters of the constituent neurons in the policy network. We applied the same DNN architecture as used in Madeka et al. (2022). The detailed training algorithm can be found in Appendix C.
@@ 4.1 Training/Evaluation Configurations
We compare our Dual Sourcing RL (DualSrc-RL) buy policy against several baselines, i.e. BaseStockHorizonTip (BSHT), Tailored Base Surge (TBS), Just-in-time RL (JIT-RL). The first two are classic operation research baselines and the JIT-RL is the RL baseline where the RL model is trained as a single-sourcing model (Madeka et al., 2022). More detailed description of those baselines are reported in Appendix A.
@@ 4.2 Main Results and Analytics
[…]
@@ 6 Conclusion and Future Work
Appendix A Baseline Methods
A.1 Improved Tailored Base Surge Policy
@@
A.2 Other Baseline Policies
@@
Appendix B Input Features
B.1 Featurization for Buying Policy
  1. 1.
    The current inventory level
  2. 2.
    Previous actions aiu that have been taken
  3. 3.
    Time series features
    (a) Historical availability corrected demand
    (b) Distance to public holidays
    (c) Historical website glance views data
  4. 4.
    Static product features
    (a) Product group
    (b) Text-based features from the product description
    (c) Brand
    (d) Volume
  5. 5.
    Economics of the product - (price, cost etc.)
  6. 6.
    Capacity costs – past costs and forecasted future costs
B.2 Featurization for Neural Coordinator
  1. 1.
    (a) Order quantities
    (b) Inventory
    (c) Availability corrected demand
    (d) Inbound
    (e) All the above, but weighted by inbound and storage volumes
  2. 2.
    Forecasted quantities for next L weeks.
    (a) Mean demand at lead time
    (b) Inventory after expected drain at lead time
    (c) All the above, but weighted by inbound and storage volumes
  3. 3.
    Other time series features
    (a) Distance to public holidays
  4. 4.
    Capacity costs (past costs and forecasted future costs)
Appendix C Training Algorithms
C.1 Training Algorithm for the Buy Policy
@@
C.2 Training Algorithm for the Neural Coordinator